4w^2 - 25 = 0 Solve by factoring and quadratic equation?
Answer by checkley79(3341)

4w^2 - 25 = 0
(2X+5)(2X-5)=0
2X+5=0
2X=-5
X=-5/2=-2.5 ANS.
2X-5=0
2X=5
X=5/2
X=2.5 ANS.